OpenStudy (anonymous):

A swimming pool has to be drained for maintenance. The pool is shaped like a cylinder with a diameter of 19 m and a depth of 1.6 m . If the water is pumped out of the pool at the rate of 15 m to the third power per hour, how many hours does it take to empty the pool? Use the value 3.14 for π , and round your answer to the nearest hour.

OpenStudy (anonymous):

hard question

OpenStudy (aivantettet26):

Differential Calculus i presume?

OpenStudy (anonymous):

wow thank you for your help @henrry123

OpenStudy (aivantettet26):

no need to be rude to him.

OpenStudy (anonymous):

I'm not being rude!

OpenStudy (kendricklamar2014):

You gals, just focus on answering @Beautykillzz question

OpenStudy (aivantettet26):

*sarcastic. anyway, 15m to the third per hour. how do you write that down?

OpenStudy (perl):

nice computer

OpenStudy (anonymous):

what do you mean how do I write that down???

OpenStudy (anonymous):

pencil + paper = write

OpenStudy (perl):

A swimming pool has to be drained for maintenance. The pool is shaped like a cylinder with a diameter of 19 m and a depth of 1.6 m . If the water is pumped out of the pool at the rate of 15 m to the third power per hour, how many hours does it take to empty the pool? Use the value 3.14 for π , and round your answer to the nearest hour. V = pi*r^2*h , here r= 19/2 , h = 1.6

OpenStudy (anonymous):

Thank you @perl !!

OpenStudy (perl):

so first calculate the volume

OpenStudy (perl):

V = 3.14 * (19/2)^2 * 1.6 = 453.416

OpenStudy (kendricklamar2014):

With what @perl gave u: = 80.4 m^3

OpenStudy (perl):

now it tells us that the rate at which it is draining is 15 m^3 / hour, so lets use d= r*t 453.416 = 15 * t

OpenStudy (kendricklamar2014):

So now u do: 80.4 / 11 = ? Hours

OpenStudy (anonymous):

so the answer would 80 hours???

OpenStudy (anonymous):

oh okay thank you so much! both of you!!

OpenStudy (perl):

I got t = 30.227 hours

OpenStudy (anonymous):

so the answer is 7 hours??

OpenStudy (kendricklamar2014):

Yes ^

OpenStudy (perl):

30 hours, rounded to the nearest hour

OpenStudy (anonymous):

thank you @KendrickLamar2014 and @perl !

OpenStudy (anonymous):

so that would be 30 hours then??

OpenStudy (perl):

yes

OpenStudy (kendricklamar2014):

Look at this: @Beautykillzz http://openstudy.com/study#/updates/4f5f972fe4b0602be43999b0

OpenStudy (radar):

@KendrickLamar2014 Note the differences in the two diameter, it is not the same problem, forget 7

OpenStudy (radar):

@aivantettet26 The flow is a rate and it could be written as:\[15 m ^{3}/hr\] now taken for a certain time it would result in Volume of cubic meters for example, after 2 hrs, a volume of 30 cubic meters of water had been removed from the pool. This is not a calculus problem like a related rate problem.

OpenStudy (aivantettet26):

@radar I see, I was confused on how it was worded. Thank you

OpenStudy (radar):

@aivantettet26 Yes the way it was written while accurate, is sort of tricky. The computer here is an iMac, yours must be a "gaming" or graphics computer, sounds like a nice set-up.

OpenStudy (aivantettet26):

@radar well technically yeah. More like it was a do-it-yourself computer rig